OpenStudy (anonymous):

which of the following lines are perpendicular to the line y=2/3x+6? (a) y=2/3x-2 (b) y=-3/2x-4 (c) y=2/3y+1 (d) y=-2/3x+3

OpenStudy (anonymous):

Perpendicular Lines means m1 * m1 = -1

OpenStudy (anonymous):

y = mx + c m = slope

OpenStudy (anonymous):

Hope nw u can Solve

OpenStudy (anonymous):

I need this explained better

OpenStudy (anonymous):

ans b y=2/3x+6 y=mx+6 m1=2/3 m2=-3/2 m1*m2=-1

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome